OpenStudy (anonymous):

Given the functions f(x) = x+2 and g(x)= x^2, what values of x result in f(g(x)) = g(f(x))

OpenStudy (anonymous):

x^2+2 = x^2+4x+2

OpenStudy (anonymous):

correct, can you solve that?

OpenStudy (anonymous):

would i solve each individually?

OpenStudy (anonymous):

I think there's a mistake above g(f(x)) = (x+2)^2 = x2 + 4x + 4, you had above +2 instead of +4

OpenStudy (anonymous):

do i need to set one side euqal to zero? or just factor out those polynomials

OpenStudy (anonymous):

if you group the x's on one side and the numbers on the other side, what do u get?

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

@ganeshie8 can I consult your expertise with one? I also get x = -1/2 just getting a second opinion...

OpenStudy (sidsiddhartha):

yes x=-1/2 seems correct

ganeshie8 (ganeshie8):

looks good to me :)

OpenStudy (anonymous):

thanks @ganeshie8 U DA MAN:)

OpenStudy (anonymous):

they both equal 2.25 :D, finally done, thanks guys

OpenStudy (anonymous):

yw